# Interactive Inference: A Neuromorphic Theory of Human-Computer Interaction

### v3

@@ Abstract.
Neuromorphic HCI is a new theoretical approach to designing better UX inspired by the neurophysiology of the brain. Here, we apply the neuroscientific theory of Active Inference to HCI, postulating that users perform Bayesian inference on progress and goal distributions to predict their next action (Interactive Inference). We show how Bayesian surprise between goal and progress distributions follows a mean square error function of the signal-to-noise ratio (SNR) of the task. However, capacity to process Bayesian surprise follows the logarithm of SNR, and errors occur when average capacity is exceeded. Our model allows the quantitative analysis of performance and error in one framework with real-time estimation of mental load. We show through mathematical theorems how three basic laws of HCI, Hick’s Law, Fitts’ Law and the Power Law fit our model. We then test the validity of the general model by empirically measuring how well it predicts human performance in a car following task. Results suggest that driver processing capacity indeed is a logarithmic function of the SNR of the distance to a lead car. This positive result provides initial evidence that Interactive Interference can work as a new theoretical underpinning for HCI, deserving further exploration.
human-computer interaction, theory, interactive inference, bayes theory
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXbooktitle: ACM Transactions on Computer-Human Interactionisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing HCI theory, concepts and models
1. Introduction
@@
There have been many efforts to develop psychological theories of human behaviour as it pertains to user interaction. Frameworks such as GOMS (Card et al., 1983) and ACT-R (Anderson et al., 1997) have been successful in allowing predictions on time performance when provided with a model user and a model interface. Similarly, models of human error (Reason, 1990) have been used to predict error as it relates to the design of interactive systems. However, there has not been a model that has comprehensively explained the relationship between human performance and error in a way that is generalizable. We believe developing such a model is important because it is the key to measuring cognitive load in real time. Current methods for measuring cognitive load fall short because they either rely on a questionnaire that can be difficult to administer during tasks (e.g., the NASA TLX (Hart, 1988)), or on unreliable physiological metrics (e.g., Heart Rate Variability, Electrodermal Activity or Pupillometry).
@@
In this paper, we propose a theory of human performance as it relates to error that allows direct measurement of performance by defining the signal-to-noise ratio (SNR) of a task. It incorporates the semantics of a task by modeling goals as predictions using Bayesian statistics. We set out to develop a unified approach called Interactive Inference that encompasses both performance and error in one information-theoretical framework. We will demonstrate through mathematical derivation that this unified approach elegantly explains existing laws of Human-Computer Interaction (HCI). It also allows for wider comparisons of information processing capacity and error between otherwise incompatible tasks. With this paper, we are advocating a more neuroscientific approach to design. The theory proposed in this work is one of the first neuromorphic contributions to HCI: an HCI that uses neuroscientific evidence of how the user’s brain learns and acts to understand how to create better user experiences.
1.1. Contributions
This paper makes the following contributions:
@@ 1.1. Contributions
    Interactive Inference model: We introduce a unified, neuromorphic framework called Interactive Inference that integrates human performance and error using information theoretical constructs. The model uses Bayesian statistics to interpret user goals as predictions, providing a new way to measure performance by quantifying the Bayesian surprise processed per signal-to-noise ratio of tasks.
@@
    Cognitive load measurement: We claim our model enables real-time estimation of cognitive load without relying on traditional self-reported questionnaires or unreliable physiological metrics. By modeling task complexity externally, it offers a more practical approach for evaluating user experience during interactions.
@@
    Derivation of HCI Laws: Through known mathematical theorems and logic, we show how foundational principles of HCI—Hick’s Law, Fitts’ Law, and the Power Law of Practice—emerge naturally from our entropy-based framework, supporting a more neuroscientifically grounded approach to interface design.
@@
    Empirical study: To verify whether Interactive Inference could successfully model a task that does not yet have a known model, we demonstrate through empirical study of a car following task that user performance follows the logarithm of the SNR of the distance to a lead car. This positive result provides initial evidence that Interactive Interference models can work as a new theoretical underpinning for HCI.
@@
    Research agenda: By discussing the limitations, benefits, and open research questions of this work, we invite further research to extend the theory and explore its practical implications. This includes further empirical studies that test the validity of the theory as applied to different task-modeling scenarios.
2. Background
2.1. Information Theory and Three Laws of HCI
@@
Shannon went on to define that the information capacity $C$ of the transmission of a message through an analog channel has a maximum that is the (binary) logarithm of the ratio between the signal strength $S$ and the amount of Gaussian noise $N$ in the channel, multiplied by some constant $b$, as expressed by the Shannon-Hartley Theorem:
(2) $C=b\cdot\log_{2}(\frac{S+N}{N})=b\cdot\log_{2}(\frac{S}{N}+1)$
Where $b$ is some empirically derived bandwidth parameter. This equation describes the capacity of information sent through the channel in terms of the signal-to-noise ratio. Note that since the negative sign is removed, the SNR represents the reciprocal probability of a transmission. A high signal allows for more uncertainty to be carried, as does a lower noise. This means the very nature of information is statistical, with underlying equations based on (normal, i.e., Gaussian) distributions. This is important because humans, unlike traditional computer algorithms, operate in a probabilistic environment. That is, their exact behaviour cannot be computed, however, it can be modeled using (normal) probability distributions. Using this framework, Card, Moran and Newell (Card et al., 1983) developed the Model Human Processor (MHP), a comprehensive set of heuristics for predicting human performance in simple computer tasks. Statistical elements to these models were derived from empirical observations such as those performed by Hick (Hick, 1952) and Fitts (Fitts, 1954). They demonstrated that human performance in certain tasks indeed seems to follow a similar pattern to the Shannon-Hartley Theorem: a binary logarithm of the signal-to-noise ratio in the task. E.g., Hick’s Law describes the response time $RT$ for selecting from a number of alternatives $n$, plus 1:
@@
Here, $b$ is some empirically derived constant that models the time per unit of difficulty, the log component of the task. Response time is derived by multiplying the amount of information in the task by this empirical constant. Fitts (Fitts, 1954) derived a very similar law describing the movement time $MT$ of a hand towards a target along a single dimensional trajectory $A$ with a target width of $W$ (specific equation from (MacKenzie, 1992)):
(4) $MT=b\cdot\log_{2}(\frac{A}{W}+1)$
One can think of the power of the signal in this equation as the amplitude (one-dimensional distance) to the target, and the noise as the width of the target. Again, the $b$ is a constant that is empirically derived. It describes an average time per bit of difficulty of the task that when multiplied with the difficulty gives the movement time prediction. It is important to note the difference between the Shannon-Hartley Theorem and these equations: these equations predict time. However, like Shannon-Hartley, all follow the Boltzmann equation for thermodynamic entropy that describes the statistical disorganization of a system (Boltzmann, 1866). This is no coincidence. While Boltzmann’s equation calculates the properties of energy distribution in physical materials, it uses statistical information on the state of the molecules in the material to do so. We will revisit this later in the paper.
The third law was derived from a paper by Snoddy (Snoddy, 1926) and describes the rate of learning of a model human. Time $T$ to complete a task follows a power law as a function of the number of trials of that task $x$ (Crossman, 1959):
@@
2.2. Predictive Coding and the Bayesian Brain Hypothesis
@@
One of the drawbacks of classic information theoretical approaches to human performance modeling is that classic information theory does not describe the presence of memory of messages transmitted through a channel, let alone proactive anticipation of those messages. According to Shannon, a bit of information is simply the negative logarithm of the probability of a message (Equation 1). Only if a message is less probable, does it contain more information. The semantics of how the human interprets that message is not modeled, nor whether the human already knows the message. This is problematic because knowing the message, or anticipating the next message, reduces its information content to zero, because this renders the uncertainty about the message zero.
Instead of framing human processing as a sequential computer processor, McClelland and Rumelhart’s parallel processing model (McClelland and Rumelhart, 1981) described perception as an interaction of top-down (cognitive/semantic) and bottom-up (sensory) elements. Rao and Ballard (Rao and Ballard, 1999) further elucidated this approach by suggesting that the brain builds up statistical knowledge of its environment, i.e., incoming messages, over time, and that this model is used generatively, i.e., top-down, to predict the subsequent messages. The error between the prediction and the messages would lead to an update of the prediction.
@@
Figure 1. Bayes Theorem. Distributions of the Likelihood P(o—s) of observations, in green, given a Prior P(s) that describes the current state, in blue. In black is the calculated Posterior P(s—o) that describes the chance of the state given the observation.
[…]
Distributions of the Likelihood P(o—s) of observations, in green, given a Prior P(s) that describes the current state, in blue. In black is the calculated Posterior P(s—o) that describes the chance of the state given the observation.
Prediction is convenient for more than one reason: not only does it allow humans to plan actions in their environment based on relevant information, it serves to compress the information content of their environment by processing only those messages that are surprising: i.e., not, or insufficiently predicted. Predictive coding theory explains how the human brain can work so efficiently and explains the mechanisms behind attention and the scheduling of cognitive resources based on the amount of surprise in their predictions (Baldi and Itti, 2010).
According to Rao and Ballard (Rao and Ballard, 1999), Bayesian inference (Bayes, 1763) is a natural candidate for a generative statistical model. Within this context, Bayesian inference estimates the probability that a particular state $s$ (representing a meaning or cause) exists, given an observation $o$. This is called the posterior probability $P(s|o)$, i.e., the chance of state $s$ after, or given observation $o$. It does so by multiplying two probabilities: the Likelihood $P(o|s)$ of the observation $o$ given state $s$, and the $Prior$ probability $P(s)$ of state $s$ before to the observation. The result is subsequently normalized by dividing by the overall probability of the observation $P(o)$ (see Figure 1):
@@
2.3. Entropy and Active Inference
@@
Friston et al. (Friston et al., 2006) further improved upon predictive coding theory by suggesting a mathematical framework for calculating the amount of Bayesian surprise (i.e., relative entropy) processed by the brain. In their framework, called Active Inference, learning and action are unified under a single principle: both use the difference between prior and (an estimate of) the posterior (i.e., the Bayesian surprise) to make decisions. Minimizing the difference informs when to learn or when to take action. When learning, the difference is mainly minimized by updating the prior to the posterior, thus updating statistical knowledge in the brain. When taking action, the difference is mainly minimized by altering the world, thereby updating the posterior to be closer to the prior knowledge of what the state of the world should be. Since we are concerned with the relative entropy between prior and posterior, the direction of updates does not principally matter.
2.3.1. Relation to Physical Entropy
To better understand the concept of relative entropy in information systems, let us first examine the concept of entropy in the physical world. Physical entropy describes the disorganization in the environment that results from most interactions between particles being random, by nature. It is governed by the second law of thermodynamics (Clausius, 1879): In a closed system, entropy (disorganization) tends to increase toward a state of thermodynamic equilibrium, but not decrease. All energy, such as heat, eventually becomes evenly distributed. Engines make use of this relative change in entropy to perform work (Clausius, 1879). Random interactions explain the natural tendency of heat to dissipate from your morning coffee, but also why groups of objects tend towards disorganization, an important realization which we will use later on. Any living organism must avoid the natural tendency towards disorganization – which equals death – by reversing this process. However, to perform this work it must borrow “free energy” from a source $outside$ the closed system of Earth: the Sun. Solar power provides food that allows all organisms to stay neatly organized. Organisms must seek food while avoiding environments that may increase their entropy, such as ones that are too cold or acidic. For this purpose, they evolved sensorimotor systems along with neural circuitry that allow them to observe, learn and act on their environment. This means the core purpose of the human sensorimotor system and cognition is to find ways to minimize the human’s physical entropy. It does so by making correct predictions about its next state, i.e., by reducing Bayesian surprise, i.e., relative $information$ entropy. Note that the parallels between reducing physical entropy and information entropy are more than skin deep. Landauer (Landauer, 1961) showed how information is exchanged for heat, which has been proven empirically (Peterson et al., 2016). This means that the equations governing information theory and the statistical properties of objects or particles are, in fact, closely related.
2.3.2. Action Mirrors Learning
One of the hypotheses of Active Inference therefore is that the approximation of Bayesian surprise, the difference in information entropy between the prior and posterior, is analogous to that of the free energy that allows organisms to maintain physical entropy. We call this approximation of Bayesian surprise free ‘information’ energy. Minimizing free ‘information’ energy is thus the purpose of all learning and action. When learning, sequential Bayesian updates of the prior allow humans to reduce their subsequent Bayesian surprise. When acting, sequential Bayesian updates of movements are also aimed at reducing Bayesian surprise, by modifying the observations instead (Körding and Wolpert, 2006). When learning, the prior moves to the posterior, when acting, the posterior moves to the prior: the roles of prior and posterior are reversed. Note, however, that mathematically these are interchangeable as the multiplication in Equation 6 is commutative. When free energy or Bayesian surprise is close to zero, learning or action stops and the goal is achieved. There is one complication in the calculation of Bayesian surprise between prior and posterior: that is estimating the probability of observations that occur outside known states, the model evidence $P(o)$. For this reason, and because the number of known states in the brain is extremely large, Active Inference uses a variational rather than Naive Bayesian approach to estimating the posterior (Domingos and Pazzani, 1996). This allows approximation of the true posterior with a simpler distribution, making the problem tractable. When this approximation is good, minimizing free ‘information’ energy becomes equivalent to minimizing Bayesian surprise. Although details of variational approaches are beyond the scope of this paper, the basic idea is to convert what we do know in the Bayes formula, the probability distributions of the likelihood P(o—s) and the prior P(s), to information via a negative log transform then summating these to find a negative log likelihood function that describes their joint probability in log space:
(8) $-log_{2}P(o,s)=-log_{2}P(o|s)-log_{2}P(s)$
@@ 2.3.2. Action Mirrors Learning
Minimizing this conveniently parabolic error function using gradient descent (Courant, 1943) can be used to approximate the minimum free energy and thus Bayesian surprise in a computationally efficient way, while ignoring the information contained in the intractable model evidence $P(o)$. This process is similar to that used for back propagation in neural networks (Rumelhart et al., 1986). Unlike classical neural networks, however, the free ‘information’ energy approach allows unsupervised real-time learning because it has an inherent reward or loss function: the minimization of Bayesian surprise does not require ground truths. In terms of the user’s experience: the reward is inherent when a system performs the function that the user expects it to perform after some input, yielding minimal Bayesian surprise and no need for further learning or action.
3. The concept of “task” as a reduction in relative entropy
@@
Active Inference implies that the core concept of HCI, the task, can be elegantly modeled as the minimizing of Bayesian surprise between prior knowledge of a desired outcome (the prior or goal) and the posterior (the progress towards a goal given current observations). This application of Active Inference to tasks we call Interactive Inference.
3.1. Reducing Physical Relative Entropy
@@
So we have a goal distribution, and a progress distribution. Let us first examine how, when we apply this concept to tasks in the real world, it becomes evident that they can be modeled by a reduction in the physical entropy of the user’s environment.
@@
For the sake of argument, consider how active inference and the corresponding reduction of entropy can be interpreted in a seemingly simple everyday task: organizing socks in a drawer. Simply because the number of possible places where your socks might end up other than the drawer is infinitely large, random interactions make it much more likely that your socks are anywhere but inside the drawer. The task, then, is to reduce the relative disorganization in the location of your socks, or their physical entropy, relative to the location of the drawer. The goal distribution describes the probability of locations where socks might be stored across the physical space inside the drawer. The progress distribution is a histogram of the probability of locations where socks may be found in the home. To execute this task and reduce the physical entropy in the organization of the socks, we must expend physical energy (walk around and pick up socks, place in drawer). So as to not violate the second law of thermodynamics, the free energy that performs this work is photosynthetic, while waste products, such as heat from our muscles, continue increasing the physical entropy of our environment in other ways.
@@
According to Active Inference, the information used by the user’s brain must follow a similar process: one that predicts any left-over physical relative entropy between socks and drawer. The brain is merely minimizing the statistical difference in information about the location of socks and drawer that also defines their physical entropy, which is in and of itself a statistical property of objects. It means that the brain, in the process of minimizing the physical entropy must minimize the information entropy of the task. To do so, it calculates the statistical likelihood that an observation of a task object corresponds to the outcome of the task, after which it can estimate the chance that the task is complete given this likelihood. According to Friston et al., it uses Bayesian inference to perform these calculations (Friston et al., 2006).
3.2. Ten Postulates on Capacity and Error in Tasks
This leads to the first of ten postulates on Interactive Inference:
@@ 3.2. Ten Postulates on Capacity and Error in Tasks
    Entropy Reduction. The aim of any task is to minimize the entropy in the space of possible task outcomes. Conversely, any errors in the execution of the task lead to increases in entropy.   The processing that is required for the user’s brain to execute a task leads to a reduction in relative information entropy, i.e., Bayesian surprise. In our example, this reduction is the information about the distribution of socks in the household. The aim of the task is to reduce the difference between where the user predicts the socks to be and where they actually are. When executing the task, the user’s brain is computing the difference between the probability distributions of locations of socks outside the drawer relative to the probability distribution of the goal: the available locations inside the drawer. In doing so it can, e.g., estimate the time and effort required to complete the task, which leads to postulates 2 and 3:
@@
    Bayesian Surprise. The reduction in relative information entropy required for a user to process a task is defined as a reduction in Bayesian surprise. Bayesian surprise is equivalent to free energy in Active Inference theory, and to relative entropy. We will use these terms as synonymous. Bayesian surprise thus equals the amount of work that needs to be performed by the brain to perform the task, in bits. Think of it as food for thought.
@@
    Scale of Task Outcomes. Relative entropy is measured along a scale of possible task outcomes $s$. This scale can be any feature of the task space that measures success or failure using some ratio level scale. In tasks in the physical world location is often used as a scale of task outcomes. However, an example in the world of information might be a score on a test.

### v4

@@ Abstract.
Neuromorphic Human-Computer Interaction (HCI) is a theoretical approach to designing better user experiences (UX) motivated by advances in the understanding of the neurophysiology of the brain. Inspired by the neuroscientific theory of Active Inference, Interactive Inference is a first example of such approach. It offers a simplified interpretation of Active Inference that allows designers to more readily apply this theory to design and evaluation. In Interactive Inference, user behaviour is modeled as Bayesian inference on progress and goal distributions that predicts the next action. We show how the error between goal and progress distributions, or Bayesian surprise, can be modeled as a simple mean square error of the signal-to-noise ratio (SNR) of a task. The problem is that the user’s capacity to process Bayesian surprise follows the logarithm of this SNR. This means errors rise quickly once average capacity is exceeded. Our model allows the quantitative analysis of performance and error using one framework that can provide real-time estimates of the mental load in users that needs to be minimized by design. We show how three basic laws of HCI, Hick’s Law, Fitts’ Law and the Power Law can be expressed using our model. We then test the validity of the model by empirically measuring how well it predicts human performance and error in a car following task. Results suggest that driver processing capacity indeed is a logarithmic function of the SNR of the distance to a lead car. This result provides initial evidence that Interactive Interference can be useful as a new theoretical design tool.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXbooktitle: ACM Transactions on Computer-Human Interactionisbn: 978-1-4503-XXXX-X/18/06
Introduction
@@
There have been many efforts to develop psychological theories of human behaviour as it pertains to user interaction. Frameworks such as GOMS (Card et al., 1983) and ACT-R (Anderson et al., 1997) have been successful in allowing predictions on time performance when provided with a model user and a model interface. Similarly, models of human error (Reason, 1990) have been used to predict error as it relates to the design of interactive systems. However, there has not been a model that has comprehensively explained the relationship between human performance and error in a way that is generalizable. We believe developing such a model is important because it is a key to measuring cognitive load of users in real time. Current methods for measuring cognitive load fall short because they either rely on a questionnaire that can be difficult to administer during tasks (e.g., the NASA TLX (Hart, 1988)), or on physiological metrics that can be difficult to interpret (e.g., Heart Rate Variability, Electrodermal Activity or Pupillometry).
@@
In this chapter, we discuss a theory of human performance as it relates to error that allows direct measurement of performance by first defining the signal-to-noise ratio (SNR) of a task. It incorporates the semantics of a task by modeling goals as predictions using Bayesian statistics. We set out to develop an approach to modeling tasks called Interactive Inference that examines both performance and error using an information-theoretical perspective. We will show how this approach might allow for wider comparisons of information processing capacity and error of users between otherwise incompatible tasks. While inspired by Active Inference, we note that our approach simplifies many of its key components and should be viewed as a practical approach rather than a strict implementation of the theory.
Key Takeaways
@@
    Interactive Inference model: We introduce a simple neuromorphic framework called Interactive Inference that integrates human performance and error using information theoretical constructs. The model uses Bayesian statistics to interpret user goals as predictions, providing a way to measure performance by quantifying the Bayesian surprise processed per signal-to-noise ratio of tasks.
@@
    Cognitive load measurement: The model enables real-time estimation of cognitive load without relying on traditional self-reported questionnaires or physiological metrics. By modeling task complexity externally, it offers a more practical approach for evaluating user experience during interactions.
@@
    Derivation of HCI Laws: Through known mathematical theorems and logic, we show how foundational principles of HCI — Hick’s Law, Fitts’ Law, and the Power Law of Practice—emerge naturally from our framework.
@@
    Empirical study: To verify whether Interactive Inference could successfully model a task that does not yet have a known model, we demonstrate through an empirical study of a car following task that user performance follows the logarithm of the SNR of the distance to a lead car.
@@
    Research agenda: By discussing the limitations, benefits, and open research questions of this work, we invite further research to extend the theory and explore its practical limitations. This includes further empirical studies that test the validity of the theory as applied to different task-modeling scenarios.
Background
Information Theory and Three Laws of HCI
@@
Shannon went on to define that the information capacity $C$ of the transmission of a message through an analog channel has a maximum that is the (binary) logarithm of the ratio between the signal strength $S$ and the amount of Gaussian noise NN in the channel, multiplied by some constant $b$, as expressed by the Shannon-Hartley Theorem:
(2) $C=b\cdot\log_{2}\left(\frac{S+N}{N}\right)=b\cdot\log_{2}\left(\frac{S}{N}+1\right)$
Where $b$ is some empirically derived bandwidth parameter. This equation describes the capacity of information sent through the channel in terms of the signal-to-noise ratio. Note that without the negative sign, the SNR represents the reciprocal of the probability of a transmission. A high signal allows for more uncertainty to be carried, as does a lower noise. This means the very nature of information is statistical, with underlying equations based on (e.g., normal) distributions. This is important because humans, unlike traditional computer algorithms, operate in a probabilistic environment. That is, their exact behaviour cannot be computed, however, it can be modeled using (normal) probability distributions. Using this framework, Card, Moran and Newell (Card et al., 1983) developed the Model Human Processor (MHP), a comprehensive set of heuristics for predicting human performance in simple computer tasks. Statistical elements to these models were derived from empirical observations such as those performed by Hick (Hick, 1952) and Fitts (Fitts, 1954). They demonstrated that human performance in certain tasks indeed seems to follow a similar pattern to the Shannon-Hartley Theorem: a binary logarithm of the signal-to-noise ratio in the task. E.g., Hick’s Law describes the response time RTRT for selecting from a number of alternatives $n$, plus 1:
@@
Here, $b$ is some empirically derived constant that models the time per unit of difficulty, the log component of the task. Response time is derived by multiplying the amount of information in the task by this empirical constant. Fitts (Fitts, 1954) derived a very similar law describing the movement time MTMT of a hand towards a target along a single dimensional trajectory AA with a target width of $W$ (specific equation from (MacKenzie, 1992)):
(4) $MT=b\cdot\log_{2}\left(\frac{A}{W}+1\right)$
One can think of the power of the signal in this equation as the amplitude (one-dimensional distance) to the target, and the noise as the width of the target. Again, the $b$ is a constant that is empirically derived. It describes an average time per bit of difficulty of the task that when multiplied with the difficulty gives the movement time prediction. It is important to note the difference between the Shannon-Hartley Theorem and these equations: these equations predict time.
The third law was derived from Snoddy (Snoddy, 1926) and describes the rate of learning of a model human. Time $T$ to complete a task follows a power law as a function of the number of trials of that task $x$ (Crossman, 1959):
@@
Predictive Coding and the Bayesian Brain Hypothesis
@@
One of the drawbacks of classic information theoretical approaches to human performance modeling is that classic information theory does not describe the presence of memory of messages transmitted through a channel, nor proactive anticipation of those messages. According to Shannon, a bit of information is simply the negative logarithm of the probability of a message (Equation 1). Only if a message is less probable, does it contain more information. The semantics of how the human interprets that message is not modeled, nor whether the human already knows the message. This is problematic because knowing the message, or anticipating the next message, reduces its information content to zero, because this renders the uncertainty about the message zero.
@@
Figure 1. Bayes Theorem. Distributions of the Likelihood P(o—s) of observations, in green, given a Prior P(s) that describes the current state, in blue. In black is the calculated Posterior P(s—o) that describes the chance of the state given the observation.:
Instead of framing human processing as a sequential computer processor, McClelland and Rumelhart’s parallel processing model (McClelland and Rumelhart, 1981) described perception as an interaction of top-down (cognitive/semantic) and bottom-up (sensory) elements. Rao and Ballard (Rao and Ballard, 1999) further elucidated this approach by suggesting that the brain builds up statistical knowledge of its environment, i.e., incoming messages, over time, and that this model is used generatively, i.e., top-down, to predict the subsequent messages. The error between the prediction and the messages would lead to an update of the prediction. Prediction is convenient for more than one reason: not only does it allow humans to plan actions in their environment based on relevant information, it serves to compress the information content of their environment by processing only those messages that are surprising: i.e., not, or insufficiently predicted. Predictive coding theory explains how the human brain can work so efficiently and explains the mechanisms behind attention and the scheduling of cognitive resources based on the amount of surprise in their predictions (Baldi and Itti, 2010). According to Rao and Ballard (Rao and Ballard, 1999), Bayesian inference (Bayes, 1763) is a natural candidate for a generative statistical model. Within this context, Bayesian inference estimates the probability that a particular state $s$ (representing a meaning or cause) exists, given an observation $o$. This is called the posterior probability $P(s|o)$, i.e., the chance of state $s$ after, or given observation $o$. It does so by multiplying two probabilities: the likelihood $P(o|s)$ of the observation $o$ given state $s$, and the PriorPrior probability $P(s)$ of state $s$ before the observation. The result is subsequently normalized by dividing by the overall probability of the observation $P(o)$ (see Figure 1):
@@
Entropy and Active Inference
@@
Friston et al. (Friston et al., 2006) further improved upon predictive coding theory by suggesting a mathematical framework for approximating the amount of Bayesian surprise (i.e., relative entropy) processed by the brain. In their framework, called Active Inference, learning and action are unified under a single principle: both use the difference between prior and (an estimate of) the posterior to make decisions. Minimizing the difference informs when to learn or take action. When learning, the difference is mainly minimized by updating the prior to the posterior, thus updating statistical knowledge in the brain. When taking action, the difference is mainly minimized by altering the world, thereby updating the posterior to be closer to the prior knowledge of what the state of the world should be. Since we are concerned with the relative entropy between prior and posterior, the direction of updates does not principally matter.
Relation to Physical Entropy
To better understand the concept of relative entropy in information systems, let us use an analogy with the concept of entropy in the physical world. Physical entropy measures the logarithm of the number of microscopic states possible in a system (Boltzmann, 1866). Because particle interactions are mostly random, systems naturally evolve toward states with higher entropy, which can be loosely described as greater disorganization. That is simply because high-entropy states are overwhelmingly more probable. This process is governed by the second law of thermodynamics (Clausius, 1879): In a closed system, entropy (analogous to disorganization) increases over time and approaches equilibrium; it does not spontaneously decrease. Eventually, energy such as heat tends to become uniformly distributed throughout the system. Engines exploit free-energy gradients, which correspond to differences in entropy, to perform work. Random interactions thus explain the natural tendency of heat to dissipate from your morning coffee into the environment. This illustrates the second law of thermodynamics: in closed systems, the entropy, which you can loosely think of as spread, increases.
Living organisms must counteract this natural tendency towards disorganization by locally staying organized (Friston, 2012). In return, they export entropy to their local environment through waste products, including heat (Prigogine, 1977). They maintain their organization by using free energy imported from outside the system, from food that concentrates solar energy. Organisms must seek food while avoiding environments that may accelerate entropy increases. For this purpose, they evolved sensorimotor systems and neural circuitry that enable them to observe, learn, and act on their environment. In humans, these systems support homeostasis much like physical systems minimize free energy: they minimize the variational free energy in their inferences, which bounds prediction error while avoiding needlessly complicated explanations (Friston, 2012). They learn to make correct predictions about their next state by reducing the Bayesian surprise in their predictions. Bayesian surprise, then, is the relative entropy between the prior and posterior beliefs about the environment. It encodes the missing information between prediction prior to observation and after observation. The equations governing the statistical properties of particles and Bayesian surprise are thus closely related.
Action Mirrors Learning
There is one complication in the calculation of Bayesian surprise between prior and posterior: that is estimating the probability of observations, the model evidence $P(o)$. For this reason, and because the number of known states in the brain is extremely large, Active Inference uses a variational rather than an exact Bayesian approach to calculating the posterior (Beal, 2003). So, Active Inference uses a minimization of Variational Free Energy (VFE) to reduce uncertainty about its predictions in a way that is mathematically analogous to the Feynman variational principle for physical free energy minimization in statistical mechanics (Feynman, 1972). It minimizes VFE because it is tractable, and from it an approximation of Bayesian surprise is obtained if the approximate posterior is close to the true posterior. In our Interactive Inference interpretation, we try to calculate Bayesian surprise directly, as it is the measure of how much our beliefs change when we learn. When learning, sequential Bayesian updates of the prior allow humans to reduce subsequent Bayesian surprise. When acting, similar sequential Bayesian updates of movements occur (Körding and Wolpert, 2006). When learning, the prior is modeled as moving to the posterior, when acting, the posterior as moving to the prior. Once Bayesian surprise is minimal, learning or action attenuates and the goal is reached. So, Interactive Inference differs from Active Inference in that we try to avoid approximating VFE by obtaining parameters that determine exact Bayesian surprise. To briefly summarize Active Inference before we continue, VFE is computed from what is tractable in Equation 6: the likelihood p(o—s) and prior p(s), which give a negative log joint probability:
(8) $log_{2}P(o,s)=-log_{2}P(o|s)-log_{2}P(s)$
@@
If distributions are normal, this produces a conveniently parabolic error function in which gradient descent (Courant, 1943) can be used to approximate the minimum free energy to subsequently approximate Bayesian surprise in a computationally efficient way, while avoiding explicit computation of the information contained in the intractable model evidence $P(o)$. This gradient descent is similar to that used for back propagation in neural networks (Rumelhart et al., 1986). Unlike classical neural networks, however, VFE allows unsupervised learning because it has an inherent reward or loss function: the minimization of free energy, which does not require ground truths. In terms of the user’s experience: the reward is inherent when a system performs the function that the user predicted it to perform after some input. VFE is minimized, and Bayesian surprise is correspondingly reduced. In the remainder of this paper, we will focus on how we might calculate Bayesian surprise more easily and more directly using the signal-to-noise ratio of a task, while remaining within the general conceptual framework of Active Inference. We emphasize that Interactive Inference simplifies many aspects of Active Inference. Particularly, we do not implement model evidence terms or expected free energy calculations, focusing instead on a tractable framework suitable for HCI applications.
The Concept of “Task” as a Reduction in Relative Entropy
@@
Interactive Inference implies that the core concept of HCI, the task, can be elegantly modeled as the minimizing of Bayesian surprise between prior knowledge of a desired outcome (the prior or goal) and the posterior (the progress towards a goal given current observations).
Reducing Physical Relative Entropy
@@
So we have a goal distribution, and a progress distribution, the relative entropy between which the brain is trying to reduce. Let us reconsider how, when we apply this concept to tasks in the real world, this can align with a reduction in the physical entropy of the user’s environment.
@@
Let us consider a seemingly simple everyday task: organizing socks in a drawer. Because the number of possible places where your socks might end up other than the drawer is infinitely large, random interactions make it much more likely that your socks are anywhere but inside the drawer. The task, then, is to reduce the relative disorganization in the location of your socks, or their physical entropy, relative to the location of the drawer. The goal distribution describes the probability of locations where socks might be stored inside the drawer. The progress distribution describes the probability of locations where socks may be found in the home. We are going to assume both are normal distributions. To execute this task and reduce the physical entropy in the organization of the socks, we must expend physical energy (walk around and pick up socks, place in drawer), i.e., perform work.
@@
According to Interactive Inference, the information used by the user’s brain follows a similar process: one that predicts any left-over physical relative entropy between socks and drawer. The brain is merely minimizing the statistical difference in information about the location of socks and drawer that also defines their relative physical entropy, which is in and of itself a statistical property of objects. It means that the brain, in the process of acting to minimize the relative physical entropy is also minimizing the relative information entropy of the task, the Bayesian surprise. To do so, it calculates the statistical likelihood that an observation of a task object corresponds to the outcome of the task, after which it can estimate the chance that the task is complete using Bayesian inference (Friston et al., 2006).
Ten Postulates on Capacity and Error in Tasks
This leads to the first of ten postulates of Interactive Inference:
@@
    Entropy Reduction. The aim of any task is to minimize the relative entropy in the space of possible task outcomes. Conversely, any errors in the execution of the task lead to increases in relative entropy.   The processing that is required for the user’s brain to execute a task leads to a reduction in relative information entropy, i.e., Bayesian surprise. In our example, this reduction is the information about the distribution of socks in the household. The aim of the task is to reduce the difference between where the user predicts the socks to be and where they actually are. When executing the task, the user’s brain is computing the difference between the probability distribution of locations of socks outside the drawer relative to the probability distribution of the goal: the available locations inside the drawer. In doing so it can, e.g., estimate the time and effort required to complete the task, which leads to postulates 2 and 3:
@@
    Bayesian Surprise. The reduction in relative information entropy required for a user to process a task is the reduction in Bayesian surprise. Bayesian surprise thus equals the amount of work that needs to be performed by the brain to perform the task, in bits.
@@
    Scale of Task Outcomes. Relative entropy is measured along a scale of possible task outcomes $s$. This scale can be any feature of the task space that measures success or failure using some ratio level scale. In tasks in the physical world relative location (i.e., distance) is often used as a scale of task outcomes. However, an example in the world of information might be a test score.